OpenStudy (anonymous):

Use the impulse-momentum theorem to find how long a stone falling straight down takes to increase its speed from 5.1 m/s to 10.6 m/s ?

OpenStudy (anonymous):

What have you done so far, bro?

OpenStudy (anonymous):

Taking down as + Force = mg Impulse = change in momentum = (10.6-5.1)m = 5.5m Impulse = Force x Time => Time = Impulse/force = 5.5/g You're welcome, feather

OpenStudy (anonymous):

I have no idea how to go upon this problem. Thank you for your help!

OpenStudy (anonymous):

I figured it out. What I did was... (10.6m/s-5.1m/s)= 5.5m/s (5.5m/s)/(9.8m/s^2)= 0.56 sec Why feather?